# Paramanu-Ayn: An Efficient Novel Generative and Instruction-tuned Language Model for Indian Legal Case Documents

### v1

@@ 6 Results and Analyses
Table 1 reports the validation perplexity of our pretrained legal models and also report the model FLOP utilization (MFU) metric for model pretraining. Table 2 displays the CPU inference speed in tokens/sec for Paramanu-Ayn pretrained models.
@@
Table 3 shows the GPT-3.5-Turbo (ChatGPT) evaluation of Paramanu-Ayn-instruct model’s responses to legal instructions related to various drafting clauses, modifications, and legal contracts across various industries on different metrics such as clarity, relevance, completeness, and legal reasoning. For GPT-3.5-Turbo evaluation, we queried ChatGPT with the respective instruction, input and our model responses from the test set to evaluate the model responses on clarity, relevance, completeness, and legal reasoning in a scale of 10. Despite Paramanu-Ayn not being pretrained on legal books, legal contracts of various types, but only exclusively on Supreme Court Case documents and with limited instruction-tuning on legal contracts, legal clauses and modifications instructions, score of 6.77/10 on clarity, 7.75/10 on relevance, 7.50/10 on completeness, and 7.75/10 on legal reasoning could be achieved. In our humble opinion, this shows the general learning ability of our models from limited instruction tuning. Table 4 evaluates the same but for only various legal instructions related to Supreme Court Cases in India. Paramanu-Ayn achieves scores of 8.00/10 on clarity, 8.89/10 on relevance, 8.22/10 on completeness, and 8.89/10 on legal reasoning. This improvement in the scores leverages the importance of relevant pretraining corpus of Supreme Court case documents.
Model Clarity Relevance Completeness Legal Reasoning
Paramanu-Ayn (8192) 6.75 7.75 7.5 7.75
Table 3: GPT-3.5-Turbo evaluation of Paramanu-Ayn-instruct responses on scale of 10 to legal instructions of drafting legal clauses and various types of legal contracts. The pretrained model was not pretrained on legal books, legal contracts.
Model Clarity Relevance Completeness Legal Reasoning
Paramanu-Ayn (8192) 8 8.888 8.222 8.889
@@
Table 4: GPT-3.5-Turbo evaluation of Paramanu-Ayn-instruct responses on scale of 10 to legal instructions related to Supreme Court cases documents.
Section B shows that our model can vividly discuss the potential legislative responses to an outcome of a Supreme Court case. Section C can identify key legal issues and summarize a given Supreme Court case in details. Section D shows that Paramanu-Ayn-instruct has the ability to draft legal argument appealing the decision of a given Supreme Court case. Section E shows that the model can also identify the legal precedents for a given case under consideration, and section F shows the model can also generate hypothetical parliamentary questions. This shows strong legal reasoning capabilities of our model. Section G shows that the model can identify potential policy changes that can be advocated for a given Supreme Court case. Section H shows the ability of the model to suggest legal strategy with explanations and constitutional reasoning for a client seeking to avoid similar legal issue for a case under consideration. Section I shows that the model can also draft hypothetical legal notice based on the facts of a Supreme Court case with reference to legal laws and constitution of India. Section J shows that the ability of Paramanu-Ayn to draft hypothetical legal petition based on a given Supreme Court case in India. Section K shows the ability of Paramanu-Ayn to draft hypothetical legal advice for a client based on the outcome of a case provided. Section L shows the ability of the model to draft legal contract generation regarding student enrollment at an academic institution, and work schedule contract generation for housing tenant. Section M shows that the model can draft work schedule legal tenant contract. Section N shows that the model can draft legal contract clause for the interest rate provision for Credit Agreement in Banking sector. Section O shows that the model can also draft financial and security clause for employment contract in detail starting from salary to employee financial benefits as mentioned in the contract including vacation days. Section P shows the model can also draft legal contract clause related to Intellectual Property rights. Section Q shows the ability of the model to draft Non-Disclosure Agreement (NDA) clause. Section R shows the model’s ability to drat legal tenant contract. All these demonstrate that Paramanu-Ayn, an exclusive pretrained legal model from scratch, which is not even considered as large in today’s Gen AI, research has the ability of powerful legal reasoning and executing different abilities of legal tasks which can be used as a legal Assistant tool for lawyers and corporate for their day to day legal address.
Section S shows the model’s response to some questions regarding the constitutional rights and legal jurisdiction in India. All these various demonstrates the powerful emerging abilities of Paramanu-Ayn-instruct despite not being pretrained on legal books, and legal contracts during pretraining, just instruction-tuning with limited examples, the pretrained Paramanu-Ayn can be instructed to draft legal clauses, legal contracts of various types, modify legal clauses. These demonstration implies that we typically do not have to pretrain model from scratch on lots of data for the model to generalize to unseen domain knowledge but a little of domain instruction tuning aids the pretrained model to gain domain knowledge through limited instruction tuning.
@@ 7 Conclusions
In this paper, we presented an exclusive Indian legal Auto Regressive Decoder based generative language legal model, Paramanu-Ayn, pretrained from scratch exclusively only on Supreme Court cases in India, Constitution of India, and Indian Penal Code for a context size of 8192 respectively. We evaluated our pretrained models on validation perplexity. We also instruction-tuned our legal model, i.e, Paramanu-Ayn-instruct on 10,763 legal instructions curated from publicly available datasets. We also reported various legal tasks capabilities of our models. Due to limited resources, we could not perform legal expert (human) evaluation on the response of our legal models to specific instructions and input. However, we evaluated the responses of prompts for instruction-tuned model, i.e, Paramanu-Ayn-instruct by GPT-3.5-Turbo on clarity, relevance, completeness, and legal reasoning metrics in a scale of 10. Paramanu-Ayn-instruct has scored above 8/10 in respective clarity, relevance, completeness, and legal reasoning metrics for legal instructions related to Supreme Court case documents. This makes us certain that our model has powerful legal reasoning and domain knowledge related to past Supreme Court cases, constitution of India, and Indian penal code and capabilities leveraging exclusive pretraining and tokenization on legal specific specilaised vocabulary.
We are also working on to develop Indian legal benchmark as the existing legal benchmarks are biased and made exclusively for Western legal laws, legislation, and penal code which is different from country specific legal jurisdiction, and legislation. In future, we would like to incorporate differential privacy (Yin and Habernal, 2022) during tokenization and pretraining on legal corpus, and continue making country specific legal language models and extend our current work for all district court cases across 28 states and 8 Union Territories in India to enable billions of people in India to leverage generative language models for their legal query and enable legal processing efficiently using AI technology.
@@
We instruction-tune our pretrained legal models on 10,763 instructions probably generated from Open AI GPT-3.5-Turbo or GPT-4, after removing the duplicates. Therefore, our instruction-tuned models may generate grammatically incoherent, and non factual legal text generation related to Indian legal jurisdiction and constitution of India as we could not verify the instruction dataset at our end. Due to limited resources, we could not measure the degree of hallucination of our legal generative language models and also could not perform legal expert human evaluation of the responses of our legal models. We also did not annonymize the publicly available legal cases during tokenization and pretraining. We also have not developed any guardrails or preprocessing of input prompts and the outputs shown in the paper are the direct responses from the model without any rule based post processing fact-checkers or toxicity classifier. It may happen our models may generate biased opinions, factually incorrect information as we kept the temperature value to 1.0 during text generation to allow the models to be highly creative.
@@ Appendix I Legal Notice
2. Whether the Supreme Court can review the appointment of a Deputy Chief Commissioner in a similar manner from another legal issue?”. Whether the Council for Scientific purposes, which has been declared invalid by the Supreme Court, should be considered a ’State’ under Article 12?“ [Your Name]”[Date]”, [Email Address][Date]”, [Law Firm’s Name]”, [Your Address]”, [Email Address]”, [Date]”, [Law Firm] Date]”, [Date]”, [Permit Address]”, [State the petitioner’s Name]”, [Specify the cause of action, etc., [Specify the cause of action, etc.”

### v2

@@ 10 Results and Analyses
Model Perplexity MFU Inference Speed
Paramanu-Ayn 1.4582 41.3476 42.4575 tokens/sec
Table 2: Perplexity, MFU, and CPU inference speed metrics (AMD epyc 7252 8-core processor, RAM: 128GB RAM)
10.1 Perplexity and Inference Speed
@@
Table 2 reports the validation perplexity of our pretrained legal models, the model FLOP utilization (MFU) metric for model pretraining, and CPU inference speed in tokens/sec for Paramanu-Ayn pretrained model.
Our validation perplexity is 1.4582 which is a sign that the legal pretrained model has been trained well. Moreover, we reach MFU of 41.3476 at bf16 mixed precision training using our in-house implementation of optimized CUDA kernels for distributed training framework. This also indicates to our training efficiency.
10.2 Legal Case Judgement Prediction with Explanation
Model Macro-F1 Accuracy (%)
Gemini Pro 0.4908 50.81
LLaMa-2 7B 0.3772 50.25
Paramanu-Ayn 0.5037 52.00
Table 3: Zero-shot evaluation on PredEX case judgement prediction with explanation task dataset (Nigam et al., 2024)
Length ROUGE-1 ROUGE-L METEOR BERTScore BLEU
1024 0.2458 0.1145 0.1548 0.4795 0.1000
2048 0.2472 0.1121 0.1974 0.5287 0.2260
4096 0.2055 0.0943 0.2080 0.5142 0.1767
5000 0.2387 0.1072 0.2400 0.4848 0.1529
6000 0.1835 0.0854 0.2060 0.5048 0.1453
8192 0.1635 0.0768 0.1964 0.4795 0.2035
Table 4: Zero-shot evaluation of Paramanu-Ayn 97M on Abstractive Supreme Court Case Summarization Test Dataset across varying summary lengths, i.e., without training on the training set. The decoding used top-p sampling as $top\_p$=0.9, and temperature was set to 1e-2. On average, the average length of test case document is 28,000 and the golden reference abstractive summary is of around 5,200.
Model ROUGE-1 ROUGE-L BLEU METEOR BERTScore
Gemini Pro 0.2484 0.1531 0.0171 0.1141 0.5759
Mixtral 7B 0.2998 0.2910 0.3022 0.3723 0.4469
Llama-3 8B 0.2937 0.1821 0.0249 0.1338 0.6204
Paramanu-Ayn 0.2387 0.1072 0.1529 0.2400 0.4848
Table 5: Zero-shot evaluation of Paramanu-Ayn 97M and LLMs on Abstractive Supreme Court Case Summarization Test Dataset, i.e., without training on the training set. All models were asked to generate summary length of 5000 tokens. The decoding used top-p sampling as $top\_p$=0.9, and temperature was set to 1e-2. On average, the average length of test case document is 28,000 and the golden reference abstractive summary is of around 5,200.
Task Dataset Accuracy (Type) Accuracy (Value)
ARC-challenge Normalized 26.45
HellaSwag Normalized 27.18
MMLU Raw 25.30
MMLU-STEM Raw 24.10
MMLU-Formal-Logic Raw 33.33
MMLU-Jurisprudence Raw 24.07
MMLU-Professional Law Raw 23.66
MMLU-International-law Raw 19.83
MMLU-Econometrics Raw 32.46
TruthfulQA-mc2 Raw 51.81
WinoGrande Raw 50.28
PIQA Raw 52.18
SciQ Normalized 40.60
COPA Raw 54.00
WIC Raw 48.75
WSC Raw 62.50
QNLI Raw 50.92
WNLI Raw 59.15
MathQA Raw 20.10
AGIEVAL-LSAT-AR Normalized 22.61
AGIEVAL-SAT-Math Raw 27.27
AGIEVAL-MATH Raw 2.90
AGIEVAL-AQuA-RAT Raw 24.02
LogiQA Normalized 23.33
Table 6: Zero-shot evaluation on various benchmarks of Paramanu-Ayn 97M.
Figure 2: GPU Utilization (%) during pretraining
Model Size AQuA-RAT SAT-Math MATH LSAT-AR LSAT-LR
MPT 7B 27.6 23.6 3.0 18.7 21.2
MPT 30B 28.0 23.9 23.9 35.1
Falcon 7B 21.7 26.4 2.3 16.1 17.3
Falcon 40B 18.5 32.7 5.5 19.6 40.2
Paramanu-Ayn (ours) 97M 24.02 27.27 2.90 22.61 20.20
Llama-1 7B 18.9 22.3 2.9 26.1 19.2
Llama-1 13B 20.1 29.5 3.9 22.2 31.6
Llama-1 33B 18.9 35.0 7.1 18.7 48.0
Llama-1 65B 23.6 41.8 10.6 23.9 56.7
Llama-2 7B 23.2 28.2 2.5 23.9 22.4
LLama-2 13B 21.7 27.3 3.9 23.0 41.0
Llama-2 34B 19.3 32.7 6.24 21.3 47.5
Llama-2 70B 23.2 41.8 13.5 25.7 70.2
Table 7: Comparison of Paramanu-Ayn 97M on AGIEVAL (Zhong et al., 2024) v/s LLM. Scores of LLM are reported from (Touvron et al., 2023). Llama was evaluated on 3-5 shot on AGIEVAL but Paramanu-Ayn 97M was evaluated at zero-shot on AGIEVAL.
We used the following zero-shot inference prompt for case judgement prediction with explanation:
“Analyze the case proceeding and predict whether the appeal/petition will be accepted (1) or rejected (0).
Subsequently provide an explanation behind this prediction with important textual evidence from the case.
### Input: case_proceeding: $<case\_pro>$
### Response: ”
Table 3 compares Paramanu-Ayn with LLMs on the case prediction with explanation task. It outperformed both LLMs significantly in both the test accuracy and macro-F1 scores despite being much smaller in size by 80 times compared to 7B LLMs and was trained for only 185 A100 hours. We also would like to point out that for domain adaptation, if a tiny generative language model is pretrained from scratch sufficiently enough, can become very competitive or even better given our findings is much more cost efficient, environmentally friendly and faster in inference by multiple order of magnitude without any optimization due to the reduced size in model parameters compared to LLMs, and reduced carbon footprint than continual pretraining of LLMs or training adapters for domain adaptation and merging with LLMs which increases the latency and inference cost drastically even one might argue on the training cost of adapters for domain adaptation.
10.3 Legal Case Abstractive Summarization
We used the following zero-shot inference prompt for case summarization:
“You are a legal assistant and your job is to summarize the underneath case proceeding given in a most concise manner while being safe.
Your summary must have the same meaning and not include false information. Make sure you do not use any external knowledge other than what is provided to you.
Your final output must only be the summarized text.
### Case: $<case>$
### Summary: ”
Table 5 shows the comparison of automatic summarization evaluation metrics of zero shot summary generation of Paramanu-Ayn versus LLMs at a summary generation length of 5000 tokens. It can be seen that Paramanu-Ayn outperforms most of the models for most of the metrics. Despite being trained only on case documents, the model has relatively performed well and even better than much larger LLMs which are trained on all kinds of variety data. This means that preprocessed legal case documents are high quality pretraining data that results in models acquiring different capabilities.
10.3.1 Impact of Length of Abstractive Summary Generation
Table 4 shows the automatic summarization evaluation of zero shot summary generation of Paramanu-Ayn 97M across varying length of abstractive summary generation. We observe that Paramanu-Ayn scored the highest BERTScore of 0.5287 at summary generation length of 2048 tokens. ROUGE-1 and ROUGE-L scores are highest for the 2048-token summary, which reflects the best match in terms of unigram overlap and sequence length with the reference summary. The METEOR and BLEU scores are also strong, suggesting a better balance between precision, recall, and n-gram overlap compared to shorter summaries.
While longer summaries tend to perform well on some metrics (e.g., BLEU and METEOR), they show a decline in ROUGE-1 and ROUGE-L scores. This suggests that extremely long summaries may lose coherence or relevance compared to the reference summary, despite capturing more information. The METEOR score improves with longer summaries up to 5000 tokens, indicating that the model’s precision and recall enhance with the length. However, very long summaries (8192 tokens) show a slight decline, suggesting a possible trade-off between completeness and accuracy. The BLEU score increases with summary length, peaking at 2048 tokens and remaining relatively high for longer summaries. This reflects better n-gram overlap but also suggests that excessively long summaries might introduce redundancy or irrelevant content.
10.4 Common-Sense Reasoning and Mathematics
Table 6 reports the zero-shot evaluation of Paramanu-Ayn across various commonsense, causal, complex reasoning tasks of natural language understanding. Paramanu-Ayn performs better than the random baseline of 25% on MMLU, i.e, a performance higher than GPT-3 models of size 13 billion parameters on zero-shot as reported in (Hendrycks et al., 2021a). Paramanu-Ayn 97M scored an impressive score of 26.45 on ARC-challenge benchmark, 25.3 on MMLU, 51.81 on TruthfulQA-mc2, 24.02 on AGIEVAL-AQuA-RAT, 27.27 on AGIEVAL-SAT-Math via zero-shot evaluation despite being exclusively pretrained on only legal case documents and having only 97 million parameters. On the mathematical benchmark, AGIEVAL-AQuA-RAT benchmark, Paramanu-Ayn 97M outperformed various models: Falcon 7B by 2.32%, Falcon 40B by 5.12% and Llama-1 7B by 5.32%, Llama-1 13B by 3.92%, Llama-1 33B by 5.32%, Llama-2 13B by 2.32%, Llama-2 34B by 4.72% points respectively. On AGIEVAL-SAT-Math, Paramanu-Ayn outperformed Falcon 7B and Llama-1 7B by 4.97%. For abstract reasoning on AGIEVAL-LSAT-AR, Paramanu-Ayn 97M outperformed Falcon 7B by 6.51%, Falcon 40B by 3.01%, Llama-1 33B by 3.91%, and Llama-2 34B by 1.31% points. The scores for Llama-1, Llama-2, and Falcon models are taken from the Llama-2 paper (Touvron et al., 2023), which used 3-5 shot evaluations for AGIEVAL. In contrast, our model was evaluated using at zero-shot setting.
TruthfulQA assesses the model’s ability to provide accurate answers, emphasizing its grasp of factual information and its capacity to avoid falsehoods (Lin et al., 2022). MMLU measures the model’s broad knowledge across various subjects including humanities, science, technology, engineering, and management (Hendrycks et al., 2021a). ARC-Challenge tests the model’s ability in complex reasoning with scientific questions (Clark et al., 2018). PIQA assesses the model’s knowledge of everyday physical processes, requiring understanding of physical commonsense (Bisk et al., 2019). COPA (The Choice of Plausible Alternatives) is a causal reasoning task for English. Given a premise and two alternatives, the task is to select the alternative that more plausibly is either the cause or effect of the premise (Gordon et al., 2012).
Paramanu-Ayn performed relatively well, achieved 62.5%on on the Winograd Schema Challenge (WSC), indicating a strong grasp of contextual and commonsense reasoning tasks. It scored 51.81% on Truthfulqa-MC2 showing reasonable accuracy in evaluating the truthfulness of statements, demonstrating effective handling of complex language understanding. It scored 52.18% on PIQA suggesting good problem-solving skills in commonsense reasoning, particularly for questions requiring everyday knowledge. Paramanu-Ayn demonstrated good performance on causal reasoning task in English; it scored 54% in COPA in zero-shot setting.
10.5 GPT-3.5-Turbo Evaluation
@@ 10.4 Common-Sense Reasoning and Mathematics
Table 8 shows the GPT-3.5-Turbo (ChatGPT) evaluation of Paramanu-Ayn-instruct model’s responses to legal instructions related to various drafting clauses, modifications, and legal contracts across various industries on different metrics such as clarity, relevance, completeness, and legal reasoning. For GPT-3.5-Turbo evaluation, we queried ChatGPT with the respective instruction, input and our model responses from the test set to evaluate the model responses on clarity, relevance, completeness, and legal reasoning in a scale of 10. Despite Paramanu-Ayn not being pretrained on legal books and legal contracts of various types, but only exclusively on Supreme Court Case documents and with limited instruction-tuning on legal contracts, legal clauses and modifications instructions, scores (out of 10) of 6.77 on clarity, 7.75 on relevance, 7.50 on completeness, and 7.75 on legal reasoning could be achieved. In our humble opinion, this shows the general learning ability of our models from limited instruction tuning. Table 8 evaluates the same but for only various legal instructions related to Supreme Court Cases in India. It shows significant improvement, with all scores of 8 and above. This improvement in the scores leverages the importance of pretraining with relevant corpus.
Task Clarity Relevance Completeness Legal Reasoning
Legal clauses and contracts 6.75 7.75 7.50 7.75
Supreme Court of India (SCI) 8.00 8.88 8.22 8.89
@@
Table 8: GPT-3.5-Turbo evaluation of Paramanu-Ayn-instruct responses on instructions related to tasks (out of 10). The evaluation was performed in February’24.
10.6 Case Studies
We performed 18 kinds of legal tasks for instruction tuning. They are listed with examples in Supplementary Material.
@@ 11 Conclusions
In this paper, we proposed an alternative approach for domain generalization, instead of doing continual pre-traning of LLMs or training adapters with LLMs on respective domain such as legal for domain adaptation or generalization, we pretrained a tiny generative language model from scratch for only 185 A100 hours of training. We presented an exclusive Indian legal auto-regressive decoder based generative language legal model, Paramanu-Ayn, pretrained from scratch exclusively only on Supreme Court cases in India, Constitution of India, and Indian Penal Code for a context size of 8192. We evaluated Paramanu-Ayn on zero-shot setting for two main legal tasks: case judgement prediction with explanation, and abstractive case summarization. Our pretrained legal model outperformed much larger LLMs trained on legal data on most of the metrics. We found out that pretraining a competitive domain specific tiny language model is more cost efficient and faster in inference by multiple order of magnitude without any optimization due to the reduced size in model parameters. If a tiny language model trained sufficiently enough, then pretraining such models for domain adaptation is much more cost efficient with low latency in inference, environmental friendly due to very less carbon footprint generation and yet very competitive to LLMs and much cost efficient, faster than continual pretraining or merging adapters to LLM which in turn increases the model parameters and increases the inference cost and latency, is indeed a new exciting research direction for domain adaptation. Furthermore, we evaluated our model on zero-shot setting across popular LLM commonsense and quantitative benchmarks for a holistic evaluation despite the model being pretrained exclusively on legal court cases documents. We also instruction-tuned our legal model, i.e, Paramanu-Ayn-instruct on 10,763 legal instructions curated from publicly available datasets, and demonstrated various legal task capabilities.
In future, we would like to extend our work for district-level court cases, which are available in Indian languages.
@@
We instruction-tune our pretrained legal models on 10,763 instructions generated from Open AI GPT-3.5-Turbo, after removing the duplicates. Therefore, our instruction-tuned models may generate grammatically incoherent, and non-factual legal text generation related to Indian legal jurisdiction and constitution of India as we could not verify the instruction dataset at our end. Due to limited resources, we could not measure the degree of hallucination of our legal generative language models and also could not perform legal expert human evaluation of the responses of our legal models. We also did not anonymize the publicly available legal cases during tokenization and pretraining. We also have not developed any guardrails or preprocessing of input prompts and the outputs shown in the paper are the direct responses from the model without any rule-based post processing fact-checker or toxicity classifier. It may happen that our models may generate biased opinions, factually incorrect information as we kept the temperature value to 1.0 during text generation to allow the models to be highly creative.
@@ Appendix I Legal Notice
2. Whether the Supreme Court can review the appointment of a Deputy Chief Commissioner in a similar manner from another legal issue?”. Whether the Council for Scientific purposes, which has been declared invalid by the Supreme Court, should be considered a ’State’ under Article 12¿‘ [Your Name]”[Date]”, [Email Address][Date]”, [Law Firm’s Name]”, [Your Address]”, [Email Address]”, [Date]”, [Law Firm] Date]”, [Date]”, [Permit Address]”, [State the petitioner’s Name]”, [Specify the cause of action, etc., [Specify the cause of action, etc.”